You estimate that Valley, Inc. will increase its dividend at 3.7 percent for the foreseeable future. The firm recently distributed a $3.68 dividend and you estimate the required return on the stock is 11.9 percent. Calculate the stock's current market value.
Answer #1

Ans $ 46.54

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 3.81616 / (11.9%- 3.7%)
P0 = 46.54
D1 = D0* (1 + g)
D1 = 3.68* (1 + 3.7%)
D1 = 3.81616